# spacing=nonfrench

### v1

@@ 3 Background
To initiate an SSH connection, both peers exchange a version banner. The Binary Packet Protocol (see below) is used from the third message on but without encryption and authentication. In the KexInit messages, nonces and ordered lists of algorithms are exchanged: One list for key exchange, one for server signatures, and two (one per direction) each for encryption, MAC, and compression. For each list, the negotiated algorithm is the first client algorithm in that list which is also offered by the server.
In the KexDhInit and KexDhReply messages, a finite-field Diffie-Hellman key exchange is performed. SSH also supports elliptic curves (ECDH) and hybrid schemes with post-quantum cryptography (PQC) as alternatives. The server authenticates itself with a digital signature as part of the handshake. The signature is computed over the contents of the previously exchanged messages, in a specified order.
@@
In contrast to TLS, SSH uses only a selection from the handshake transcript for authentication. The hash value computed from this selection is called exchange hash $H$, defined as:
$H=\text{HASH}(V_{C}~{}||~{}V_{S}~{}||~{}I_{C}~{}||~{}I_{S}~{}||~{}K_{S}~{}||~{}X~{}||~{}K)$
where HASH is the hash function of the negotiated key exchange, $V_{C}$ and $V_{S}$ are the version banners of the client and server, $I_{C}$ and $I_{S}$ are the KexInit messages, $K_{S}$ is the server’s public host key, and $K$ the shared secret derived from the key exchange. The value of $X$ depends on the key exchange and contains a composition of negotiated parameters (if any) and the ephemeral public keys of the key exchange [29, Sec. 8]. Each field includes a length field defined by the encoding.
@@
Each sequence number is stored as a 4-byte unsigned integer initialized to zero upon connection. After a binary packet has been sent or received, the corresponding sequence number $\mathsf{Snd}$ or $\mathsf{Rcv}$ is incremented by one. Sequence numbers are never reset for a connection but roll over to 0 after $12^{32}-$ 1. To avoid replay attacks, rekeying must occur at least once every $2^{32}$ packets [33, Sec. 6.1].
@@
Figure 3: CBC-EaM in SSH (diagram based on [2, Fig. 1]).
@@
The BPP is used to encrypt and authenticate messages. First, a message is prefixed by a 4-byte message length and a 1-byte padding length. Then, at least 4 bytes of padding are added to the message so that the total plaintext length is a multiple of the block size or 8, whatever is larger. On the secure channel, the packet is encrypted by the cipher mode, and a MAC is added. The details depend on the chosen cipher mode, which makes use of an implicit initialization vector $\mathrm{IV}_{\mathrm{KDF}}$ derived from the session key:
@@
CBC-EaM [29] is part of the original SSH specification. Here, the MAC is computed over the implicit sequence number and the packet plaintext (Figure 3). The IV of the first packet is $\mathrm{IV}_{\mathrm{KDF}}$, and IV chaining is used (i.e., the IV of packet $i$ is the last ciphertext block of packet $i-1$).
@@
CBC-EtM [32] was added to OpenSSH in 2012. Here, the packet length is not encrypted to allow checking the MAC before decryption. The MAC is computed over the sequence number, the unencrypted packet length, and the ciphertext of the packet payload. The IVs are handled as with CBC-EaM.
@@
CTR [33] mode was proposed by Bellare, Kohno, and Namprempre [3] as a countermeasure to attacks on CBC with IV chaining. Here, $\mathrm{IV}_{\mathrm{KDF}}$ is used as the initial counter value and incremented after encrypting a plaintext block. CTR can be used with EaM or EtM, with identical implications for the length field and MAC computation as above.
@@
GCM [20] mode was specified by the NSA for Suite B-compliant SSH implementations [19]. Here, ciphertext integrity is part of the encryption scheme. The length field is not encrypted (solely authenticated) to allow verification of the authentication tag before returning any plaintext. Internally, GCM uses an invocation counter that is initialized to $\mathrm{IV}_{\mathrm{KDF}}$ and incremented by one for each message. The sequence number is not used but is always offset by a constant from the invocation counter.
@@
ChaCha20-Poly1305 [30] was added to OpenSSH in 2013, inspired by a similar proposal for TLS by Langley and Chang [23, 24]. Here, two different encryption keys are derived, one for the length field and one for the packet payload. This prevents the length field from being used as a decryption oracle for the payload. The MAC is computed over the concatenation of the two ciphertexts. Internally, the AEAD construction uses the sequence number as IV for each packet.
@@
We note that the SSH specification says that the length field is encrypted [29, Sec. 6], and that the sequence number is used for integrity checks [29, Sec. 6.4]. This is only true for CBC-EaM, CTR-EatM, and ChaCha20-Poly1305. The modes CBC-EtM, CTR-EtM, and GCM do not encrypt the length field, and GCM also does not use the sequence number.
@@ 4 Breaking SSH Channel Integrity
As detailed in Section 3, the exchange hash signed by the server during the handshake only authenticates some parts of the handshake transcript, while other parts are left unauthenticated. This allows an attacker to inject some messages into the handshake, which cannot affect the key exchange but can affect the implicit sequence numbers of the peers.
SSH Does Not Reset Sequence Numbers at the Beginning if the Secure Channel
In SSH, sequence numbers are only incremented and never reset to 0, even when the encryption state changes. This allows an attacker to manipulate the sequence number counters in the secure channel before encryption and authentication keys are activated.
@@
In IPsec/IKE, only a portion of the handshake transcript is signed, but unlike SSH, sequence numbers are reset to 0 when encryption and MAC keys are activated. In TLS, Finished messages are exchanged at the beginning of the secure channel to verify the integrity of the complete handshake transcript, and sequence numbers are reset to 0 after every ChangeCipherSpec message.
@@ 4.1 Sequence Number Manipulation
In this section, we show how a MitM attacker can arbitrarily increase the receive sequence numbers $\mathsf{C.Rcv}$ and $\mathsf{S.Rcv}$ in the client and the server. This will be the basis for our prefix truncation attack and its applications, allowing the attacker to compensate for messages deleted from the secure channel.
Technique RcvIncrease (5(a))
During the handshake, a MitM attacker can increase $\mathsf{C.Rcv}$ (resp. $\mathsf{S.Rcv}$) by $N$, while not changing any other sequence number, by sending $N$ Ignore messages to the client (resp. server).
The correctness is evident from the fact that the SSH standard requires for Ignore that “All implementations MUST understand (and ignore) this message at any time.” [29, Sec. 11.2]. The intended purpose of this message is to protect against traffic analysis, so it is considered a security feature, although there is no benefit from it during the handshake phase. We note that the attacker may also use any other message type that does not generate a response.
In addition, we found that an attacker can set the sequence numbers to arbitrary values by abusing the rollover after $2^{32}$ messages. These advanced techniques require that the implementation allows handshakes with many messages, a large amount of data, and a long operating time. As we do not require these advanced techniques for our attacks, a description can be found in Appendix A.
@@
We verified that, as expected, this technique works with OpenSSH 9.4p1 and 9.5p1, Dropbear 2022.83, PuTTY 0.79, AsyncSSH 2.13.2, and libssh 0.10.5.
@@ 4.2 Prefix Truncation Attack on the BPP
We first analyze this attack with regard to handshake authentication and sequence numbers. As the key exchange does not protect the handshake transcript from inserting Ignore messages (Section 3), handshake authentication is not broken. Before the first step, we have $\mathsf{C.Rcv}=\mathsf{C.Snd}$. After the first step, we have $\mathsf{C.Rcv}=\mathsf{S.Snd}+1$, but during the handshake this manipulation is not detected. After the second step, we have $\mathsf{C.Rcv}=\mathsf{S.Snd}$, and sequence numbers are back in sync.
It remains to be shown that the attacker can delete the message from the channel, which requires knowing its length, and that its deletion does not affect the MAC verification and decryption output for the following messages. This analysis depends on the encryption mode and will be given in Section 4.3. Here, we conclude by describing the general attack.
@@
In a single attack, the attacker can generally delete an arbitrary number of $N_{S}$ initial messages sent from the server and $N_{C}$ initial messages sent from the client. This is straightforward: Instead of inserting one Ignore message to the client before NewKeys, the attacker inserts $N_{S}$ such messages to the client and $N_{C}$ to the server. Consequently, instead of deleting the first message from the server, the attacker deletes $N_{S}$ initial messages from the server and the $N_{C}$ initial messages from the client.
@@
4.3 Analyis of Encryption Modes
In this section we first discuss how an attacker can determine the byte-length of SSH messages. Then we analyze which encryption modes are vulnerable to our attacks, and which can be exploited. An encryption mode is vulnerable if, after prefix truncation, all following packets on the secure channel are decrypted, i.e. an AEAD mode does not generate the distinguished symbol Invalid or an authenticated encryption mode successfully verifies the MAC. Note that we allow decryption to a different plaintext for probabilistic attacks. To capture this, we define that an encryption mode is exploitable for an attack, if the message stream after decryption is well-formed and supports that attack.
Determining the Byte-Length of Messages
To successfully delete packets from the secure channel, the attacker has to know their length. For some encryption modes (CBC-EtM, CTR-EtM, GCM), the length is sent unencrypted so the attacker can observe it. For others (CBC-EaM, ChaCha20-Poly1305), the length is encrypted, and the attacker has to guess the length, either from known plaintext based on the used SSH implementations or from network side channels such as TCP segment sizes. For the following analysis, we assume that the attacker always knows the message lengths.
4.3.1 Not Vulnerable
@@
GCM [20] mode does not use the implicit sequence number. Instead, it uses an invocation counter to derive encryption and MAC keys, initialized to $\mathrm{IV}_{\mathrm{KDF}}$ and incremented after each message. The authors justify this by stating that the resulting nonce is always a fixed offset from the sequence number. By deviating from the SSH standard, GCM stops our attack, as the attacker cannot manipulate the invocation counter during the handshake.
@@
CBC uses IV chaining, and CTR uses a key stream. When the attacker deletes any prefix of the ciphertext in either mode, the first ciphertext block received will be decrypted as pseudo-random. Because EaM computes the MAC over the plaintext, MAC verification will fail with probability close to 1, and our attack is stopped.
4.3.2 Vulnerable And Perfectly Exploitable

### v2

@@ 3 Background
To initiate an SSH connection, both peers exchange a version banner. The Binary Packet Protocol (see below) is used from the third message on but without encryption and authentication. In the KexInit messages, nonces and ordered lists of algorithms are exchanged: One list for key exchange, one for server signatures, and two (one per direction) each for encryption, MAC, and compression. For each list, the negotiated algorithm is the first algorithm in the client’s list, which is also offered by the server.
In the KexDhInit and KexDhReply messages, a finite-field Diffie-Hellman key exchange is performed. SSH also supports elliptic curves (ECDH) and hybrid schemes with post-quantum cryptography (PQC) as alternatives. The server authenticates itself with a digital signature as part of the handshake. The signature is computed over the contents of the previously exchanged messages in a specified order.
@@
In contrast to TLS, SSH uses only a selection from the handshake transcript for authentication. The hash value computed from this selection is called exchange hash $H$, defined as
$H=\text{HASH}(V_{C}~{}||~{}V_{S}~{}||~{}I_{C}~{}||~{}I_{S}~{}||~{}K_{S}~{}||~{}X~{}||~{}K),$
where HASH is the hash function of the negotiated key exchange, $V_{C}$ and $V_{S}$ are the version banners of the client and server, $I_{C}$ and $I_{S}$ are the KexInit messages, $K_{S}$ is the server’s public host key, and $K$ is the shared secret derived from the key exchange. The value of $X$ depends on the key exchange and contains a composition of negotiated parameters (if any) and the ephemeral public keys of the key exchange [33, Sec. 8]. Each field includes a length field defined by the encoding.
@@
Each sequence number is stored as a 4-byte unsigned integer initialized to zero upon connection. After a binary packet has been sent or received, the corresponding sequence number $\mathsf{Snd}$ or $\mathsf{Rcv}$ is incremented by one. Sequence numbers are never reset for a connection but roll over to 0 after $12^{32}-$ 1. As sequence numbers are responsible for protecting against replay attacks, rekeying must occur at least once every $2^{32}$ packets [37, Sec. 6.1].
@@
(a) Encrypt-and-MAC [33] (diagram based on [2, Fig. 1]). The packet length is encrypted as part of the packet’s plaintext. The sequence number is part of the input used to compute the MAC.
(b) Encrypt-then-MAC [36, Sec. 1.5] (diagram based on [41, Slide 37]). The packet length is not encrypted but authenticated as part of the MAC’s input. The sequence number is used similarly to Encrypt-and-MAC.
(c) Galois/Counter Mode [22]. The packet length is not encrypted but authenticated as additional authenticated data (AAD). The sequence number is not used directly but is replaced with an invocation counter of constant offset.
(d) ChaCha20-Poly1305 [34] (diagram based on [41, Slide 39]). The packet length is encrypted using a different key and authenticated as part of the Poly1305 input. The sequence number is encoded as an unsigned 64-bit integer to match the nonce length of ChaCha20.
Figure 3: Commonly used authenticated encryption schemes in the BPP of SSH.
@@
The BPP is used to encrypt and authenticate messages. First, a message is prefixed by a 4-byte message length and a 1-byte padding length. Then, at least 4 bytes of padding are added to the message so that the total length is a multiple of the block size or 8, whatever is larger. On the secure channel, the packet is encrypted by the cipher mode, and a MAC is added. The details depend on the authenticated encryption scheme, which uses an implicit initialization vector $\mathrm{IV}_{\mathrm{KDF}}$ derived from the session key.
@@
CBC-EaM [33] (3(a)) is part of the original SSH specification. The MAC is computed over the implicit sequence number and the packet plaintext. The IV of the first packet is $\mathrm{IV}_{\mathrm{KDF}}$, and IV chaining is used (i.e., the IV of packet $i$ is the last ciphertext block of packet $i-1$).
@@
CBC-EtM [36] (3(b)) was added to OpenSSH in 2012. Here, the packet length is not encrypted to allow checking the MAC before decryption. The MAC is computed over the sequence number, the unencrypted packet length, and the ciphertext. The IVs are handled as with CBC-EaM.
@@
CTR [37] mode was proposed by Bellare, Kohno, and Namprempre [4] as a countermeasure to attacks on CBC with IV chaining. $\mathrm{IV}_{\mathrm{KDF}}$ is used as the initial counter value and incremented after encrypting a plaintext block. CTR can be used with EaM or EtM, with identical implications for the length field and MAC computation as above.
@@
GCM [22] (3(c)) mode was specified by the NSA for Suite B-compliant SSH implementations [21]. Here, ciphertext integrity is part of the cipher mode. The length field is not encrypted (solely authenticated) to allow verification of the authentication tag before returning any plaintext. Internally, GCM uses a 12-byte nonce that is initialized to $\mathrm{IV}_{\mathrm{KDF}}$. The nonce is split into a 4-byte fixed value and an 8-byte invocation counter that is incremented by one for each message. The sequence number is not used but is always offset by a constant from the invocation counter.
@@
ChaCha20-Poly1305 [34] (3(d)) was added to OpenSSH in 2013, inspired by a similar proposal for TLS by Langley and Chang [26, 27]. Here, two different encryption keys for the length field and the packet payload are derived, so the length field cannot be used as a decryption oracle for the payload. The MAC is computed over the concatenation of the two ciphertexts. Internally, the AEAD construction uses the sequence number as a nonce for each packet.
@@
We note that the SSH specification says that the length field is encrypted [33, Sec. 6] and that the sequence number is used for integrity checks [33, Sec. 6.4]. This is only true for CBC-EaM, CTR-EatM, and ChaCha20-Poly1305. The modes CBC-EtM, CTR-EtM, and GCM do not encrypt the length field, and GCM also does not use the sequence number.
@@ 4 Breaking SSH Channel Integrity
As detailed in Section 3, the exchange hash signed by the server during the handshake only authenticates some parts of the handshake transcript, while other parts are left unauthenticated. This allows an attacker to inject messages into the handshake, which cannot affect the key exchange but does affect the implicit sequence numbers of the peers.
SSH Does Not Reset Sequence Numbers at the Beginning of the Secure Channel
In SSH, sequence numbers are only incremented and never reset to 0, even when the encryption key changes. This allows an attacker to manipulate the sequence number counters in the secure channel before encryption and authentication keys are activated.
@@
In IPsec/IKE, only a portion of the handshake transcript is signed, but unlike SSH, sequence numbers are reset to 0 when encryption and MAC keys are activated. In TLS, Finished messages are exchanged at the beginning of the secure channel to verify the integrity of the complete handshake transcript, and sequence numbers are reset to 0 after installing new keys. The Noise Protocol Framework fully secures the handshake transcript and uses a nonce as a sequence counter that is initialized to 0 after the handshake.
@@ 4.1 Sequence Number Manipulation
In this section, we first show how a MitM attacker can arbitrarily increase the receive sequence numbers $\mathsf{C.Rcv}$ and $\mathsf{S.Rcv}$ in the client and the server during the handshake. This will be the basis for our prefix truncation attack and its applications, allowing the attacker to compensate for messages deleted from the secure channel.
Technique RcvIncrease (4(a))
A MitM attacker can increase $\mathsf{C.Rcv}$ (resp. $\mathsf{S.Rcv}$) by $N$ while not changing any other sequence number by sending $N$ Ignore messages to the client (resp. server).
The correctness is evident from the fact that the SSH standard requires for Ignore that “All implementations MUST understand (and ignore) this message at any time.” [33, Sec. 11.2]. The intended purpose of this message is to protect against traffic analysis, so it is considered a security feature, although there is no benefit from it during the handshake phase. We note that the attacker may also use any other message type that does not generate a response.
Other Modifications of Sequence Numbers
In addition, we found that an attacker can set the sequence numbers to arbitrary values by using the rollover after $2^{32}$ messages during the handshake. These advanced techniques require that the implementation allows handshakes with many messages, a large amount of data, and a long operating time. We also require a message that generates a response message but is otherwise ignored. Conveniently, the SSH standard requires this for all messages with unrecognized message IDs [38, Sec. 11.4]. Let Unknown be a message with an unrecognized message ID.
Technique RcvDecrease (4(b))
A MitM attacker can decrease $\mathsf{C.Rcv}$ (resp. $\mathsf{S.Rcv}$) by $N$ while not changing any other sequence number by sending $2^{32}-N$ Ignore messages to the client (resp. server).
Technique SndIncrease (4(c)) and SndDecrease (4(d))
A MitM can increase $\mathsf{C.Snd}$ (resp. $\mathsf{S.Snd}$) by $N$ while not changing any other sequence number by sending $N$ Unknown and $2^{32}-N$ Ignore messages to the client (resp. server) and deleting all generated Unimplemented messages. Conversely, a MitM can decrease $\mathsf{C.Snd}$ (resp. $\mathsf{S.Snd}$) by $N$ while not changing any other sequence number by sending $2^{32}-N$ Unknown and $N$ Ignore messages to the client (resp. server) and deleting all generated Unimplemented messages.
Here, the total data transfer required is $\approx 69\mathrm{GB}$ for SndIncrease and twice as much ($\approx 138\mathrm{GB}$) for SndDecrease. Again, these techniques may fail on implementations that have timeouts or restrict the amount of data or number of messages exchanged during the handshake.
@@
We verified all techniques successfully against PuTTY 0.79. Additionally, our experiments show that OpenSSH 9.5p1 recognizes a rollover of sequence numbers and terminates the connection, thus not affected by any technique but RcvIncrease. AsyncSSH 2.13.2 and libssh 0.10.5 allow for RcvIncrease but terminate the connection due to handshake timeouts before any advanced technique concludes. Dropbear 2022.83 disconnects on Unknown messages instead of responding with Unimplemented but allows $\mathsf{Rcv}$ to roll over, therefore being affected by RcvIncrease and RcvDecrease only.
(a) RcvIncrease
(b) RcvDecrease
(c) SndIncrease
(d) SndDecrease
@@ 4.2 Prefix Truncation Attack on the BPP
We first analyze this attack with regard to handshake authentication and sequence numbers. As the key exchange does not protect the handshake transcript from inserting Ignore messages (Section 3), handshake authentication is not broken. Before the first step, we have $\mathsf{C.Rcv}=\mathsf{S.Snd}$. After the first step, we have $\mathsf{C.Rcv}=\mathsf{S.Snd}+1$, but this manipulation is not detected during the handshake. After the second step, we have $\mathsf{C.Rcv}=\mathsf{S.Snd}$, and sequence numbers are back in sync.
It remains to be shown that the attacker can delete the message from the channel, which requires knowledge about the message’s length, and that its deletion does not affect the MAC verification and decryption output for the following messages. Both aspects require careful analysis with respect to the used encryption mode, which will be given in Section 4.3 and Section 4.4. Here, we conclude by describing a straightforward generalization of the single message attack.
@@
In a single attack, the attacker can generally delete an arbitrary number of $N_{S}$ initial messages sent from the server and $N_{C}$ initial messages sent from the client. This is straightforward: Instead of inserting one Ignore message to the client before NewKeys, the attacker inserts $N_{S}$ such messages to the client and $N_{C}$ to the server. Consequently, instead of deleting the first message from the server, the attacker deletes $N_{S}$ initial messages from the server and $N_{C}$ initial messages from the client.
@@
4.3 Determining the Byte-Length of Messages
To successfully delete packets from the secure channel, the attacker has to know their length. This is inherently true for encryption modes that do not encrypt the packet length field (any EtM mode, GCM). In the case of an encryption mode with an encrypted packet length field (any EaM mode, ChaCha20-Poly1305), the attacker may employ different strategies to determine the packet’s length. One such strategy is to utilize knowledge about the plaintext if the length of the first few messages inside the secure channel is either fixed (for example, ServiceAccept) or can be measured within a single connection ahead of time (for example, ExtInfo). This approach was used for all attacks described here. More advanced strategies may exploit TCP segment sizes and timings, as well as the message order of the SSH protocol. For example, an attacker may delay all encrypted traffic by the server until after the client’s ServiceRequest message has been processed to determine the length of the ExtInfo message. Here, we assume that the attacker always knows the lengths.
4.4 Analysis of Encryption Modes
Authenticated Encryption Mode Specification Enc. State Dec. State Affected Exploitable Ref.
Encrypt-and-MAC CBC [33] ($IV$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Snd}}}$) ($IV$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Rcv}}}$) Section 4.4.1
CTR [33, 37] ($ctr$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Snd}}}$) ($ctr$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Rcv}}}$) Section 4.4.1
Encrypt-then-MAC CBC [33, 36] ($IV$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Snd}}}$) ($IV$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Rcv}}}$) Section 4.4.3
CTR [37, 36] ($ctr$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Snd}}}$) ($ctr$, ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Rcv}}}$) Section 4.4.3
GCM [22] $ctr_{Invocation}$ $ctr_{Invocation}$ Section 4.4.1
ChaCha20-Poly1305 [34] ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Snd}}}$ ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\mathsf{Rcv}}}$ Section 4.4.2
Table 1: Authenticated encryption modes, corresponding specification documents, and their exposure to prefix truncation in the BPP of SSH. The initial value of state variables printed in bold purple can be chosen by the attacker, cf. Section 4.1. Full control of either state enables perfect prefix truncation (●, ChaCha20-Poly1305). Partial control may lead to limited exploitability, depending on the inner workings of the authenticated encryption mode (◐, Encrypt-then-MAC).
In this section, we analyze which encryption modes our attacks affect and if they can be exploited in a real-world scenario (see Table 1). An encryption mode is affected if, after prefix truncation, all following packets on the secure channel are decrypted, i.e., an AEAD mode does not generate the distinguished symbol Invalid or a composed mode successfully verifies the MAC. Note that we allow decryption to a different plaintext for probabilistic attacks. To capture this, we define an encryption mode as exploitable for an attack if the message stream after decryption is well-formed and supports that attack. If the attack’s success probability is less than 1, we say the attack has limited exploitability.
4.4.1 Not Affected
@@
GCM [22] mode does not use the implicit sequence number. Instead, it uses an invocation counter, initialized to $\mathrm{IV}_{\mathrm{KDF}}$, and incremented after each message. The authors justify this by stating that the resulting nonce is always a fixed offset from the sequence number. By deviating from the SSH standard, GCM stops our attack, as the attacker cannot manipulate the invocation counter during the handshake.
@@
CBC uses IV chaining, and CTR uses a key stream. When the attacker deletes any prefix of the ciphertext in either mode, the first ciphertext block received will be decrypted as pseudorandom. Because EaM computes the MAC over the plaintext, MAC verification will fail with a probability close to 1, thwarting our attack.
4.4.2 Affected And Perfectly Exploitable